OpenStudy (anonymous):

If the average adult woman consumes approximately 2,000 kcal per day, how long would she need to run in order to utilize 25 percent of her caloric intake, given that the energy requirement for running is 42,000 J per minute?

OpenStudy (shane_b):

To utilize 25% of her daily intake calories she will need to burn:\[0.25*2000kcal=500kcal\] You're given her rate of calorie burning in J/min...so convert that to kcal/min knowing that 1kcal is about 4200J: \[\frac{42,000 \cancel{J}}{min}*\frac{1kcal}{4200\cancel{J}}=10kcal/min\] Therefore, at 10kcal/min it will take her: \[\frac{500kcal/min}{10kcal}=50min\] So it will take her 50 minutes...

OpenStudy (shane_b):

I posted that last fraction incorrectly...it should have been:\[\frac{500kcal}{10kcal/min}=50min\]

OpenStudy (anonymous):

Thank you soo much!

OpenStudy (shane_b):

np :)